4. On October 1, a gardener plants 20 bulbs. On October 2, she plants 23
bulbs. On October 3, she plants 26 bulbs. She continues in this pattern
until October 15, on which she plants the last bulbs.
a. Write an explicit formula to model the number of bulbs she plants
each day.
b. Write a recursive formula to model the number of bulbs she plants
each day.
c. How many bulbs will the gardener plant on October 15?
d. What is the total number of bulbs she plants from October 1 to
October 15, inclusive?

Answer by RAY100(1637)

a(n) =a(1) + (n-1) d,,,,where a(n) = nth entry,,n=no entries,,d = difference between entries
.
a(15) = 20 + (15-1) 3 = 20 + 42 = 62
.
recursive, a(n) = a (n-1) +d
.
sum = n(a(1) +a(n) ) /2= 15 *( 20 +62)/2 = 615
.